### v2

# GraphXAIN: Narratives to Explain Graph Neural Networks

Abstract
Graph Neural Networks (GNNs) are a powerful technique for machine learning on graph-structured data, yet they pose interpretability challenges, especially for non-expert users. Existing GNN explanation methods often yield technical outputs such as subgraphs and feature importance scores, which are not easily understood. Building on recent insights from social science and other Explainable AI (XAI) methods, we propose GraphXAIN, a natural language narrative that explains individual predictions made by GNNs. We present a model-agnostic and explainer-agnostic XAI approach that complements graph explainers by generating GraphXAINs, using Large Language Models (LLMs) and integrating graph data, individual predictions from GNNs, explanatory subgraphs, and feature importances. We define XAI Narratives and XAI Descriptions, highlighting their distinctions and emphasizing the importance of narrative principles in effective explanations. By incorporating natural language narratives, our approach supports graph practitioners and non-expert users, aligning with social science research on explainability and enhancing user understanding and trust in complex GNN models. We demonstrate GraphXAIN’s capabilities on a real-world graph dataset, illustrating how its generated narratives can aid understanding compared to traditional graph explainer outputs or other descriptive explanation methods.
-0.7cm0cm
1 Introduction
The exponential growth in the complexity of machine learning models has led to architectures reaching billions of parameters, resulting in significant improvements in their performance. [19, 29, 8, 3]. As these complex ‘black-box’ models, characterized by their high accuracy yet lack of interpretability, continue to evolve, the demand for transparency has intensified [26]. Explainable Artificial Intelligence (XAI) has emerged to address this challenge by enhancing the trustworthiness and transparency of complex decision-making processes [12].
Graph Neural Networks (GNNs) [30] have recently gained notable success and have become state-of-the-art solutions for modelling relational data characterized by nodes connected via edges [33, 35, 15]. However, the need for interpretability in GNNs remains [36, 18, 16].
Existing GNN explanation methods primarily provide explanatory subgraphs and feature importance values, as illustrated in the upper section of Figure 1, which displays an explanation generated by GNNExplainer [36] to explain the GNN model’s prediction for node 57 in the NBA dataset, based on player’s attributes, fields statistics and social connections. Interpreting the subgraph visualization alongside feature importance metrics alone can pose significant challenges for readers. Without a complementary narrative, end users must rely solely on the technical outputs of subgraphs and feature importance to interpret the model’s underlying reasoning, which can be particularly challenging given the complex nature of GNN models. Despite its limitations, GNNExplainer remains a state-of-the-art graph explainer.
Such an approach is not aligned with the comprehensibility postulate introduced by Michalski (1983) [23],
which advocates that computer-generated results should be directly interpretable in natural language. Moreover, effective explanations should enhance the alignment between the user’s mental model and AI model [21, 14]. GNNExplainer’s approach, among other graph explainers, does not adhere to these principles and exemplifies the phenomenon termed the inmates running the asylum [25], where solutions are technical and cater primarily to experts, neglecting the needs of non-technical users who prefer explanations in natural language [4]. Martens et al. (2023) [20] proposed XAIstories, a framework that employs Large Language Models (LLMs) to generate narratives explaining AI model’s predictions based on SHAP [19] and counterfactual (CF) [22, 34] explanations. In their survey among data scientists and lay users, over 90% of the general audience found these narratives convincing, while 83% of the surveyed data scientists indicated they would likely use XAIstories to communicate explanations to non-expert audiences. By integrating narrative communication as a compliment to the XAI explanations, we can present the model’s decision-making process in a manner that aligns with human cognitive preferences, enhancing explanations beyond mere descriptions and effectively illustrating cause-and-effect relationships [4].
To the best of our knowledge, no previous research proposes the general framework for natural language XAI Narratives for Graph Neural Network models. In this article, we introduce the first method to automatically generate GraphXAINs, natural language narratives, for explaining GNN models. By complementing explanatory subgraphs and feature importances with coherent XAI Narratives, we aim to present the decision-making processes of GNNs in a more transparent and accessible way to end users. We posit that this framework will not only enhance interpretability but also facilitate more effective communication of model insights across various graph applications.
To summarise, our main contributions are as follows:
We present a novel model-agnostic and explainer-agnostic method that generates natural language XAI Narratives to enhance the explainability of prediction models over graph-structured data.
We illustrate our approach by integrating GraphXAIN with the existing graph XAI framework (GNNExplainer) for GNNs and demonstrating its explanatory abilities on a real-world dataset.
We formalise the concepts of narrative and descriptive explanations within the context of XAI, clarifying their distinctions and discussing their implications for model interpretability.
2 Related Work
2.1 Explainability in Machine Learning
Several approaches have been proposed to enhance the explainability of machine learning models across various modalities, including image data [29, 31, 32, 1, 11], tabular data [29, 19, 2], natural language [32, 19, 29], and unstructured data such as graphs [36, 18]. One of the most popular XAI methodologies are post-hoc explanations, which aim to interpret deep learning models after the training stage [26]. These methods encompass feature importance measures, visualisation techniques, and surrogate models. For instance, SHAP (SHapley Additive exPlanations) [19] and LIME (Local Interpretable Model-agnostic Explanations) [29] estimate the contribution of each feature to a particular prediction, thereby shedding light on the model’s decision-making process for specific instances.
However, despite these advancements, challenges persist in ensuring that explanations are both methodologically accurate [7] and meaningful to end users [20, 27]. Explanations must bridge the gap between technical complexity and user comprehension, necessitating a careful balance between fidelity and interpretability [4, 25]. The need for explainability methods that are understandable to non-technical users is primarily crucial in sensitive domains such as healthcare, finance, and legal systems, where understanding the deep learning model’s decision-making process is essential for trust and transparency [29, 26].
2.2 Graph Neural Networks Explainability
GNNs are increasingly used for modelling relational data in domains such as social networks, molecular structures, and knowledge graphs [30, 15, 33]. Their complex architectures, however, pose challenges for understanding and interpreting their predictions. The first method developed to address this issue is GNNExplainer [36], which provides instance-level explanations by identifying a compact subgraph and relevant node features that are most influential for a specific prediction. GNNExplainer formulates the explanation task as an optimisation problem, maximising the mutual information between the explanatory subgraph with a subset of node features, and the input graph that is subject to explain.
Among other graph explanation methods, Lucic et al. (2021) [18] introduced CF-GNNExplainer which alters the GNNExplainer to answer ’what-if’ questions using the counterfactual explanation approach. Rather than merely identifying influential features or subgraphs, CF-GNNExplainer searches for minimal perturbations to the input graph that would alter the GNN model’s prediction by edge deletion. This method demonstrates how small changes in graph structure or node features impact outcomes, enhancing understanding of the model’s decision-making process.
Although the aforementioned state-of-the-art graph explanation frameworks are methodologically sound, they merely provide users with additional graphs and node feature importance values that are not easily interpretable by end-users, thereby limiting their practical utility and compelling users to construct the explanatory narrative themselves.
Incorporating natural language into GNN explanations could bridge the gap between technical outputs and human understanding by translating complex model reasoning into accessible narratives, thereby enhancing user comprehension and trust. However, previous methods to generate them are not tailored to the popular GNNExplainer outputs and/or provide descriptions rather than narratives (see below for a discussion of the important differences). Giorgi et al. (2024) [6] addressed this issue by using LLMs to generate textual explanations for counterfactual graphs. However, their explanations lack contextual information and do not illustrate cause-and-effect relationships, resulting in descriptive communication rather than narrative explanations [4] (see Appendix B.2 for examples).
He et al. (2024) [10] used LLMs to generate natural language explanations for counterfactual graphs in the context of molecular property prediction. However, since their framework adheres to domain-specific knowledge, it cannot be considered a general method for graph model explanations. Furthermore, the explanations produced exhibit the same limitation as in Giorgi et al. (2024) [6], being more descriptive rather than narrative (see Appendix B.3 for examples).
The most comparable approach to ours is presented by Pan et al. (2024) [27], who developed TAGExplainer, a method for generating natural language explanations for Text-Attributed Graphs (TAGs). Although their explanations incorporate elements of narrative communication, their framework is limited only to TAGs. This limitation restricts the generalisability of TAGExplainer to broader graph data modelling contexts, as TAGs represent only a subset of real-world graph data (see Appendix B.4 for examples). Moreover, TAGExplainer does not provide an explanatory subgraph or feature importance values.
3 Methods
3.1 XAI Narratives and Descriptions
Research in communication theory and psychology indicates that narrative-based explanations are more accessible and memorable than descriptive methods, making them effective for conveying scientific evidence to non-expert audiences [4]. Moreover, narratives are processed more rapidly by individuals without prior knowledge and are more engaging and persuasive, thereby enhancing trust and understanding of AI models [4, 25, 20].
Further, as narrative communication relies on contextual cause-and-effect relationships, it is considerably more challenging to fragment a narrative into smaller, meaningful segments without either significantly altering the interpretation of these segments or disrupting the coherence of the original narrative [4]. Consequently, narratives are often perceived as storytelling, characterised by a coherent structure comprising an introduction, main body, and conclusion. In contrast, descriptive, context-free communication can be readily fragmented into smaller units while still effectively conveying the necessary information [4]. The XAIstories [20] framework addresses aforementioned explanation limitations by enhancing the narrative communication of SHAP and CF explanations, aligning with the research on human-AI interactions [4, 25].
Having identified the need to distinguish between narrative and descriptive explanations in the context of XAI, we propose definitions for both terms, drawing on research from the social sciences [4, 25].
Definition 1 (XAI Narrative)
A XAI Narrative provides a structured, story-like representation of a model’s prediction. Narrative explanations illustrate the relationships between key features in a context-dependent manner, providing a coherent and comprehensive understanding of how the model arrives at specific outcomes.
Definition 2 (XAI Description)
A XAI Description provides a static presentation of key features or attributes relevant to a model’s prediction, delivered in a context-free and fact-based manner.
An example description would list the most important features and neighbouring nodes. This relates closely to the data-to-text or graph-to-text [28, 5]. Figure 5 provides an XAI Description (XAID) example for the XAI solution provided by GNNExplainer shown in Figure 1. Clearly, an XAID is less valuable than an XAIN, as descriptions are less accessible and memorable than narrative communication methods, making them less effective for conveying scientific evidence to a broader audience [4, 20].
3.2 From GNN models to Natural Language Narratives
To overcome the issue of purely technical explanations of GNN models’ predictions, we propose GraphXAINs—natural language narratives for graphs. We propose the following definition of GraphXAIN, a XAI Narratives for graphs:
Definition 3 (GraphXAIN)
A GraphXAIN is a XAI Narrative tailored for graph-structured data.
Our solution involves converting subgraph structures and feature importance metrics derived from graph explainers into GraphXAINs - coherent natural language narratives. The workflow, presented in Figure 2, begins with the training of a GNN classification model on a relational graph dataset. Our approach is not limited to pure graphs; data can include associated node and edge features. Further, using the graph explainer method, we extract relevant subgraphs and feature importances. These technical outputs are then transformed and prompted into an LLM for further inference. The LLM generates a natural language narrative that explains the prediction of the target node’s label, highlighting the contributions of specific neighbouring nodes, corresponding edges, and node features in a context-dependent manner, resulting in a coherent cause-and-effect narrative presented to the end user.
Importantly, our framework is agnostic to the graph data type, graph model, and graph explainer, allowing its application across various graph scenarios. By translating technical explanations into coherent narratives, we aim to make graph model explanations more intuitive and accessible for practitioners.
4 Experiments
4.1 Dataset
We conducted our experiments on a real-world graph NBA dataset111https://www.kaggle.com/noahgift/social-power-nba, which includes players’ performance statistics from the NAB 2016-2017 season, alongside various personal attributes such as height, weight, age, and nationality. The graph was constructed by linking NBA players based on their relationships on the Twitter platform. After preprocessing, filtering out disconnected nodes, and ensuring the graph is undirected, where any connection on the Twitter platform results in a mutual edge, the final graph comprised of the total number of 400 nodes (players), 42 node features, and 21,242 edges (Twitter connections).
The node classification task involves predicting whether a player’s salary exceeds the median salary. Prior to training, the data were randomly divided into training, validation, and test sets in a 60/20/20 split.
4.2 Graph Model
We trained a two-layer Graph Convolutional Network (GCN) [15] model with 16 hidden channels. For training we used the Binary Cross-Entropy loss function and the AdamW [17] optimizer with a learning rate of 0.001 and a weight decay of 510-4. The training process continued through 1400 epochs, resulting in the GCN model achieving a 0.80 test AUC.
4.3 Graph Explainer
To obtain subgraph and feature importance explanations, we used GNNExplainer [36], the current state-of-the-art method for explaining GNN models. The GNNExplainer formulates the explanation task as an optimisation problem, aiming to maximise the mutual information between the explanatory subgraph and a subset of node features, relative to the input graph under consideration. The training process for GNNExplainer was conducted for 200 epochs, adhering to the default settings recommended for this explainer by the authors. Nevertheless, in our framework, any graph explainer may be used as long as it provides an explanatory subgraph.
4.4 Large Language Models and Prompts
XAI Narratives are derived from the GNNExplainer output, with the use of LLM (GPT-4o). Using a specifically developed prompt, we guide the LLM to produce coherent and relevant narratives that accurately reflect the graph model’s prediction. The language model is provided with a prompt containing the target node’s feature values, along with feature importance and the subgraph derived from GNNExplainer. Additionally, feature values of the nodes within the subgraph and the final GNN model prediction for the target node are included. In our visualisations, we use the seven most important features and restrict the subgraph to seven nodes, aligning with Miller’s (1956) [24] theory on cognitive limits, which suggests that seven pieces of information represent an optimal amount for receiving, processing and recalling. The full prompt schema is presented in Appendix A.
5 Results
In the following, we provide examples demonstrating various automatically generated GraphXAINs to explain the GNN model’s final prediction of the target node. It is important to emphasize that the results presented are not selectively chosen. From the graph dataset, five node indices were randomly selected. The GraphXAINs are subsequently generated for each of the selected nodes and are presented in Figure 3 and Figure 4. For the remaining GraphXAIN examples, see Appendix B.1.
The results of the GraphXAINs demonstrate the effectiveness of generating XAI narrative-based explanations from the context-free subgraphs and feature importance values. Unlike other GNN XAI descriptive outputs, which typically present static feature values without context, the outputs generated by our method provide a coherent XAI Narrative. Not only do these narratives articulate which features contributed most to the model’s prediction, but also explain how and why these features combined in a cause-and-effect manner. For instance, the narrative for a high-salary prediction highlights the player’s field and performance statistics, such as position and the number of games played, while crucially contextualising these within broader patterns of team dynamics and social interactions.
Figure 5 presents a possible XAI Description for the GNN model’s final prediction of node 57, outlining key features contributing to the target label prediction. The XAI Description is presented in a static, context-free, and non-coherent format, emphasising the fundamental distinctions between a description and a narrative. The superiority of the XAI Narrative presented at bottom of the Figure 1 over the provided XAI Description is evident. Thus GraphXAIN goes beyond the XAID as shown in Figure 5.
This approach ensures that the explanation is not limited to mere numerical and visual descriptions, but instead offers a comprehensive, story-like narrative that enables the end-user to understand the model’s decision-making process more intuitively. By conveying the reasoning behind the prediction, the narrative XAI explanation helps to bridge the gap between technical model outputs and user comprehension, which is a key purpose of effective narrative communication [4].
6 Conclusion and Future Work
In this work, we have introduced GraphXAIN, a novel solution that transforms explanatory subgraphs and feature importance metrics into coherent XAI Narrative explanations for GNN models. The proposed framework addresses a significant gap in the current landscape of Explainable AI by moving beyond technical, descriptive outputs and providing intuitive, story-like explanations that enhance user comprehension and trust. Our results demonstrate that XAI Narratives can deliver clearer insights into the decision-making processes of complex GNN models, making these processes more accessible to non-expert users.
For future work, it will be essential to test our framework on other real-world graph datasets to further validate its effectiveness across different domains. Moreover, due to the unstructured nature of graph data, LLMs may be more prone to generate hallucinations, since they struggle to interpret such type of data with the same accuracy as structured inputs [13]. This issue could potentially be mitigated by integrating the proposed method with an approach tailored for graph inputs for LLMs like for instance, the G-Retriever [9]. Additionally, there is a need to develop robust quantitative and qualitative methodologies to assess the accuracy of the generated XAI Narratives.
Data and Code Availability
The data and code used in this research are available at: https://github.com/ADMAntwerp/GraphXAIN
Acknowledgments
This research received funding from the Flemish Government under the “Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen”.
Appendix A Prompt Schema
Figure 6 presents the schema of the prompt used within the proposed framework.
Appendix B Examples of XAI Explanations in Natural Language for GNN Models
B.1 GraphXAIN
In this section, we present the remaining examples generated by our framework. Figure 7 and Figure 8 provide further insight into the interpretability of the GNN’s model prediction by presenting coherence narratives, which demonstrates the GraphXAIN’s consistency across a range of randomly sampled nodes. Each example was generated following the same methodology, with nodes randomly sampled using the random.sample(range(0, len(data.x)), 5) function and set_seed(42) to ensure reproducibility. The examples showcase the explanatory depth of GraphXAINs in capturing the reasoning behind the GNN model’s predictions.
In the truncated subgraph visualisations presented in both Figure 7 and Figure 8, GNNExplainer is instructed to reduce the subgraph to the seven most influential nodes according to edges’ weight importance. As a result, this process occasionally results in disconnected nodes within the truncated subgraph, as GNNExplainer prioritises node importance over connectivity in the simplified view. Consequently, these subgraph representations can appear more unintuitive or fragmented. A complementary natural language narrative is therefore argued to be essential to bridge these interpretative gaps, providing end users with a coherent understanding of the GNN model’s final prediction, regardless of the connectivity of the subgraph.
For example, to obtain a connected subgraph for node 379, GNNExplainer requires at least 14 nodes in this particular example. In this larger subgraph, previously disconnected nodes 7 and 263, become connected to the target node 379, revealing their substantial influence on the final prediction. However, as the subgraph expands, it becomes increasingly complex and challenging to interpret intuitively. Thus, a complementary natural language narrative is needed to provide a coherent understanding of the GNN model’s prediction. Figure 9 presents GraphXAIN’s output for a connected subgraph containing the 14 most influential nodes for the target node’s 379 prediction.
B.2 Natural Language Counterfactual Explanations for Graphs Using Large Language Models (Giorgi et al., 2024)
This section presents three examples of textual explanations proposed by Giorgi et al. (2024) [6] for counterfactual explanations of node classification using a GNN model trained on the Cora dataset. The Cora dataset is a standard citation network dataset in which nodes represent research papers, and edges denote citation relationships, with the task being to classify each paper into one of seven categories. Since the authors do not provide any output examples in their article, the examples presented below are taken from the article’s official GitHub repository. These examples were not selectively chosen but represent the first three explanations generated by the best-performing model, as reported in the paper and available in the article’s official GitHub repository222https://github.com/flaat/llm-graph-cf/blob/main/data/results/cf-gnn_Qwen2.5-14B-Instruct-GPTQ-Int4_cora_Response.json:
“The target node (node 15) was originally classified as ‘Probabilistic_Methods’ in the factual graph. However, in the counterfactual graph, the classification changes to ‘Reinforcement_Learning’. Despite the features of node 15 remaining identical, the classification shifts due to changes in the classification of its neighboring nodes. Specifically, nodes 4 and 12, which were previously classified as ‘Probabilistic_Methods’, are now reclassified as ‘Reinforcement_Learning’. This change in the neighborhood context strongly influences the classification of node 15, leading to its reclassification as ‘Reinforcement_Learning’. The only neighbor that remains consistent is node 19, but its influence is outweighed by the changes in nodes 4 and 12. Thus, the primary factor driving the change in classification is the altered classification of neighboring nodes, rather than any intrinsic change in the features of node 15 itself.” (Line 4)
“In the factual scenario, Node 3 is classified as ‘Theory’ and is connected to Nodes 0, 9, 10, and 13. In the counterfactual scenario, Node 3 is reclassified as ‘Case_Based’ and is only connected to Node 10. The change in classification is influenced by the loss of connections to Nodes 0, 9, and 13, and the retention of a connection to Node 10, which is classified as ‘Case_Based’.” (Line 8)
“In the factual scenario, Node 0 is classified as ‘Theory’ and is connected to both Nodes 1 and 4. Node 1 shares similar features with Node 0 and is also classified as ‘Theory’. However, Node 4, which is classified as ‘Case_Based’, introduces a conflicting influence. In the counterfactual scenario, the connection to Node 4 is removed, isolating Node 0 from the ‘Case_Based’ influence. As a result, Node 0 is reclassified as ‘Case_Based’, reflecting the reduced influence of ‘Theory’ nodes and the absence of conflicting ‘Case_Based’ connections. This counterfactual explanation highlights how the removal of a single connection can significantly alter the classification of a node, providing insight into the decision-making process of the graph-based model.” (Line 12)
According to Dahlstrom (2014) [4] and our proposed definitions, these examples constitute an XAI Description rather than an XAI Narrative.
B.3 Explaining Graph Neural Networks with Large Language Models: A Counterfactual Perspective for Molecular Property Prediction (He et al., 2024)
Presented below are three examples of textual explanations generated by the method proposed by He et al. (2024) [10] for counterfactual explanations in the context of molecular property prediction using a Graph Neural Network (GNN) model trained on a chemical molecule dataset. This dataset comprises molecular structures in which nodes represent atoms and edges signify chemical bonds, with the primary objective being the prediction of specific molecular properties. These three examples are the only instances provided by the authors in their article [10]:
“This molecule contains a cyclohexane ring, a dithiane ring, a ketone group, and a thiocarbonyl group, in which the ketone group may be the most influential for AIDS treatment.” (Page 3, Figure 2)
“The molecule contains hydroxylamine, cyclohexane, sulfone, and thioether functional groups, in which hydroxylamine may be the most influential for AIDS treatment.” (Page 4, Figure 4)
“This molecule contains a cyclohexane ring, a dithiane ring, a ketone group, and a hydrazine group, in which the hydrazine group may be the most influential for AIDS treatment.” (Page 6, Figure 5)
Again, the explanations presented by He et al. (2024) [10] are delivered in a context-free and fact-based manner, without illustrating underlying relationships or cause-and-effect relationships, thereby categorising them as XAI Descriptions rather than XAI Narratives.
B.4 TAGExplainer: Narrating Graph Explanations for Text-Attributed Graph Learning Models (Pan et al., 2024)
Presented below are three instances of textual explanations proposed by Pan et al. (2024) [27] for Text-Attributed Graph (TAG) learning models using their TAGExplainer method. The authors applied TAGExplainer to the Cora, DBLP, and Book-History datasets to generate natural language explanations that incorporate elements of narrative communication. These three examples are the only instances provided by the authors in their article, each corresponding to the aforementioned datasets, respectively:
“The classification of the ROOT node into the ‘Reinforcement Learning’ category can be explained as follows:
ROOT mentions reinforcement learning, learning algorithms, and hierarchical control structures indicating that it focuses on reinforcement learning methodologies and their applications. - Node-1 discusses transfer learning in reinforcement learning, highlighting how learning agents can adapt to multiple tasks.
Node-1.1 - Node-1.8 provide more context about various reinforcement learning architectures and approaches, including modular Q-learning and hierarchical reinforcement learning.
Node-2 focuses on learning in continuous domains with delayed rewards, a pertinent aspect of reinforcement learning, particularly in real-world applications.
Node-2.1 discusses generalization in reinforcement learning using neural networks, which is relevant but less directly tied to the prediction.
Node-3: This node discusses the need for hierarchies within control systems and the improvements in nested Q-learning, which are highly relevant to the classification label.
In summary, the classification of the ROOT node into ‘Reinforcement Learning’ is well-supported by the strong presence of key reinforcement learning terms within the node itself and its direct connections to Nodes 1, 2 and 3, which also focus on reinforcement learning methodologies.” (Page 10, Figure 4b)
“The classification of the ROOT node into the ‘Cloud computing’ category can be explained as follows:
ROOT: The presence of keywords such as ‘user’, ‘interfaces’, and ‘service-oriented approach’ indicates that the paper is focused on topics central to cloud computing, particularly how users interact with cloud services and the modular nature of these services.
Node-1: While Node 1 discusses ‘graphical user interfaces’ and ‘domain-specific languages’, which are relevant to cloud computing, it is slightly less directly related to the core concepts of cloud computing compared to the ROOT node.
Node-1.1: This node emphasizes ‘user interface requirements’ and ‘requirements analysis’, which are crucial for developing effective cloud applications that meet user needs.
Node-1.2: This node focuses on ‘stereotypes’ and ‘domain-specific terms’, highlighting the importance of understanding user interactions and the context in which cloud services are utilized.
In summary, the classification of the ROOT node into ‘Cloud computing’ is well-supported by the presence of key terms related to user interaction and service-oriented architectures. The direct connection to Node 1, which discusses user interfaces, further reinforces this classification, while the additional insights from Nodes 1.1 and 1.2 emphasize the importance of user-centric design in cloud computing applications.” (Page 16, Figure 6b)
“The classification of the ROOT node into the ‘Europe’ category can be explained as follows:
ROOT: The presence of keywords such as ‘cambodia’, ‘year’, and ‘translation’ indicates a context that may involve historical or cultural discussions relevant to Europe, particularly in terms of colonial history and cultural exchanges. The mention of ‘english’ and ‘french’ highlights the linguistic dimensions that are significant in European contexts.
Node-1: This node discusses Michael D. Coe, an anthropologist specializing in Southeast Asia and the Khmer civilization. While it provides historical context, the focus on Southeast Asia may dilute its direct relevance to Europe. However, the terms ‘civilizations’ and ‘ancient’ could connect to European historical interests.
Node-2: This node is more directly relevant as it discusses the destruction of Cambodia during the Nixon-Kissinger era, a significant historical event that involved European powers’ interests in Southeast Asia. The emphasis on ‘destruction’ and ‘cambodia’ alongside key historical figures suggests a critical perspective on the geopolitical dynamics involving European countries.
In summary, the classification of the ROOT node into ‘Europe’ is supported by the presence of key terms that indicate a historical and cultural context relevant to European interests, particularly through the stronger connection found in Node-2.” (Page 16, Figure 7b)
The presented examples align with the proposed XAI Narrative definition. However, the TAGExplainer approach applies only to text-attributed graphs and does not provide an explanatory subgraph or feature importance.